Circular No. 8570

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005da IN UGC 11301 Further to IAUC 8563, E. Lee and W. Li report the LOSS discovery

of an apparent supernova (mag 16.9) on unfiltered KAIT images taken on July 18.35 and 19.36 UT. SN 2005da is located at α = 18^{h}37^{m}46^{s}.89, δ = +17^{o}32′35′′.3 (equinox 2000.0), which is 108′′.4 west and 34′′.8 north of the nucleus of UGC 11301. A KAIT image taken on 2004 June 2.36 showed nothing at this position (limiting mag ∼ 19.5).

OCCULTATION BY PLUTO I (CHARON) L. A. Young, C. B. Olkin, and E. F. Young, Southwest Research Insti-

tute; and R. G. French, Wellesley College, report their observations of an occultation on July 11 of the star 2UCAC 26257135 by Pluto I (Charon). E. F. Young and K. Shoemaker observed with 0.2-s integration times using the visitor instrument PHOT (unfiltered) on the 4.2-m SOAR telescope at Cerro Pachon. C. Olkin, C. Ruhland, and L. A. Young observed via a 240-s I-band exposure trailed at 1′′.5/s with the 4-m Blanco telescope (+ Mosaic II camera) at Cerro Tololo. R. French, B. Gregory, L. A. Young, and R. Galvez observed via 0.5-s I-band exposures using the 0.9-m SMARTS telescope (+ Tek2K No. 3 CCD camera) at Cerro Tololo. All three sites report durations of 55.3 ± 0.2 s. The chord length is 1179 ± 4 km, which serves as a lower limit on the diameter from these telescopes. No obvious atmosphere was detected. Immersion and emersion times (± 0^{s}.01) at SOAR were 3^{h}36^{m}16^{s}.19 and 3^{h}37^{m}11^{s}.26 UT.

COMETS C/2005 L10, C/2005 L11, C/2005 L12, C/2005 L13 (SOHO)

K. Battams reports data concerning additional tailless Kreutz sungrazing comets found on SOHO website images (cf. IAUC 8569). C/2005 L10 was tiny and very faint. C/2005 L11 was stellar in appearance, reaching mag ∼ 6.4 at 7.0R_{⊙} on June 10.142 UT. C/2005 L12 was stellar in C3 images but very diffuse in C2 images, reaching mag ∼ 7.1 at 7.0R_{⊙} on June 12.292. C/2005 L13 was very diffuse (estimated mag 8.0).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 L10 June 9.433 5^{h}07.^{m}7 +21^{o}06′ C2  KC 2005-N75
C/2005 L11  9.596 5 10.3 +19 49 C3/2 BZ 2005-N76
C/2005 L12  12.113 5 19.2 +20 52 C3/2 TH 2005-N76
C/2005 L13  12.117 5 17.0 +21 12 C2  HS 2005-N76
2005 July 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green